OpenStudy (anonymous):

Using the domain values {-1, 0, 4}, find the range values for relation y=2x-7. Range={-1, 0, 4} Range={-9, -7, 1} Range={-5, -7, 9} None of the above

OpenStudy (anonymous):

its the second

hero (hero):

Hint: x = {-1, 0,4}

OpenStudy (anonymous):

Basically pluck in the numbers from the domain, and whatever you get for the y= will be your range.

hero (hero):

Please don't give answers @curiousman

OpenStudy (anonymous):

domain is x and range is y, so what u gotta do is substitute the x values which are the domains into the equation to solve for y, range, for all of these values :)

OpenStudy (anonymous):

why? thank you everyone!!!

OpenStudy (anonymous):

like for example, if the domain is 2, then what is the range. y= 2x - 4 y= 2(2) - 4 y = 4-4 y= 0 so then when x, the domain is 2, then y, range is 0 :) u got it?